# Topological components of surface group representations and the Signature

### v1

@@ 2.4. The case of general genus
If $\mathrm{tr}(C_{1})>2$, then we can similarly choose deformations $A_{1}(t),B_{1}(t)$ such that $\mathrm{tr}(C_{1}(t))$ decreases into the interval (2,2)(-2,2), while preserving the sign of $(C_{1}(t))_{21}$. In fact, by the path-lifting property [8, Corollary 7.8], we can construct deformations $A_{1}(t),B_{1}(t)$ such that $\mathrm{sgn}((C_{1}(t))_{21})$ remains constant throughout the deformation, where $C(t)$ is hyperbolic for $[0,1)$, and $C(1)$ is parabolic. Moreover, we can arrange that
@@
For representations satisfying $|\mathrm{T}(\phi)|<|\chi(\Sigma_{g,0})|$, the argument is the same as in Lemmas 2.20 and 2.23, as detailed in [8, Lemma 10.5]. This shows that for each such Toledo invariant value, the set of corresponding representations forms a connected component, and there are 2g32g-3 such values.
@@ 2.5. The case of $\mathrm{PSL}(2,\mathbb{R})$
Let $\pi:\mathrm{SL}(2,\mathbb{R})\to\mathrm{PSL}(2,\mathbb{R})$ denote the natural projection. This induces a map which is a covering map on its inverse of a connected component ([8, Lemma 2.2])
@@
Since $\phi^{\prime}|_{\pi_{1}(\Sigma_{g,0})}]$ and $\psi^{\prime}|_{\pi_{1}(\Sigma_{g,0})}]$ are connected in $\mathrm{Hom}(\pi_{1}(\Sigma_{g,0}),\mathrm{PSL}(2,\mathbb{R}))$ (see [8, Theorem A (i)]), and $\phi^{\prime}|_{\pi_{1}(\Sigma_{0,n})}$ and $\psi^{\prime}|_{\pi_{1}(\Sigma_{0,n})}$ are also connected, it follows that $\phi^{\prime}]$ and $\psi^{\prime}]$ are connected in $\mathrm{Ell}_{P}(\Sigma_{g,n})$. Consequently, $\phi]=[\mu(\phi)]$ and $\psi]$ lie in the same connected component of $\mathrm{Ell}_{P}(\Sigma_{g,n})$. ∎
@@ 3.2. The general case
Assume the lemma holds for all $\leq n$. For $n+1$, consider the maximal dual-tree decomposition as illustrated on the right in Figure 4. By [8, Lemma 10.1], we may assume that $D_{1},\dots,D_{n-2}$ are hyperbolic elements.
@@
By [8, Lemma 10.1], we may assume that $C_{-1}$ and $C_{1}$ are hyperbolic. If $\mathrm{T}(\phi_{1})<1$, then as in the case $n=3,g=0$, we may deform $\phi_{1}$ along a $\gamma$-path while fixing the conjugacy classes of $C_{-1}$ and $C_{0}$, so that the Toledo invariant of $\phi_{1}$ increases by 1.
@@
By [8, Lemma 10.1], we may assume that it is an interior hyperbolic representation with respect to a maximal dual tree decomposition, see also Lemma 4.10.
@@
Note that $\mathrm{Hyp}(\Sigma_{1,1})\cap\mathrm{T}^{-1}(0)$ is connected. This follows from the same argument as in [8, Lemma 10.5] and the sigma value of each representation in $\mathrm{Hyp}(\Sigma_{1,1})\cap\mathrm{T}^{-1}(0)$ is always equal to $1$ (see Table 2), since one can construct a representation $\phi_{1}\in\mathrm{Hyp}(\Sigma_{1,1})\cap\mathrm{T}^{-1}(0)$ such that $\phi_{1}(a_{1})$ or $\phi_{1}(b_{1})$ is elliptic. Furthermore, if the set $\mathrm{Hyp}(\Sigma_{0,g+n})\cap\mathrm{T}^{-1}(l)\cap\sigma^{-1}(a)$ is non-empty, by Theorem 3.3 and the fact that $\sigma^{-1}(a)$ is fixed, it is connected. By Theorem 3.3, any $\phi$ as above can be deformed into a representation of the form $\gamma\phi_{0}$, where $\gamma\in\Lambda_{g,n}$.
@@
Write $C_{3}:=\Pi_{i=1}^{g-1}\phi([a_{i},b_{i}])$ and $C_{2}:=\phi([a_{g},b_{g}])$, so that $C_{3}C_{2}C_{1}=I$. By [8, Lemma 10.1], we may assume that $C_{2}$ and $C_{3}$ are both hyperbolic. Since $\mathrm{T}(\phi)=\pm|\chi(\Sigma_{g,1})|=\pm(2g-1)$, and by the inductive assumption, we have $\sigma(C_{2})=\sigma(C_{3})=0$. Furthermore, the triple $(C_{1},C_{2},C_{3})$ forms a representation of $\Sigma_{0,3}$ with Toledo invariant zero if and only if $\sum_{i=1}^{3}\sigma(C_{i})$ is odd. Hence, $T((C_{1},C_{2},C_{3}))=\pm 1$ implies $\sigma(C_{1})=0$.
We now consider the case $n\geq 2$. Again, by [8, Lemma 10.1], we may assume that $C_{0}$, $D_{1},\dots,D_{n-2}$ are all hyperbolic (see Figure 4).
@@ 4.1.1. The case of pairs of pants
In this section, we consider the case $g=0$, that is, $\Sigma=\Sigma_{0,n}$, a genus-zero surface with $n$ boundary components.
@@
(4.8) $\begin{split}&\quad\#\{a_{1},a_{2}:\mu s=1,c\in\{0,2\}\}+\#\{a_{1},a_{2}:\mu s=-1,c=0,s+\mu a^{2}\neq 0\}\\ &=2\,\#\{a_{1},a_{2}:\mu s=1\}+2\#\{a_{1},a_{2}:\mu s=-1\}\\ &=2\times 8+2\times 8=32.\end{split}$
@@ 4.2. Interior hyperbolic representations
In [8, Lemma 10.1], Goldman proved that any boundary hyperbolic representation $\phi\in\mathrm{Hyp}(\Sigma)$ can be deformed to another boundary hyperbolic representation $\phi^{\prime}$, such that for a maximal dual tree decomposition $\Sigma=\bigcup_{i=1}^{|\chi(\Sigma)|}M_{i}$ (see [8, Section 3.8]), the restriction $\phi^{\prime}|_{\pi_{1}(M_{i})}$ is hyperbolic on $\partial M_{i}$ for each $i$. We refer to such a representation $\phi^{\prime}$ as an interior hyperbolic representation.
In this subsection, we demonstrate that any boundary parabolic representation can similarly be continuously deformed into one that is hyperbolic on every interior pair of pants.
@@
Theorem 4.6 (Goldman [8, Proposition 4.1, Theorem 4.3, Corollary 4.5]).
@@
Let $\Sigma=\Sigma_{g,n}$ be a surface of genus $g$ and $n$ boundaries. We assume that $\chi(\Sigma)\leq-1$. Let $\Sigma=\bigcup_{i=1}^{|\chi(\Sigma)|}M_{i}$ be a maximal dual tree decomposition. Let $\mathcal{R}_{hp}(\Sigma)$ be the set of all representations $\phi\in\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SL}(2,\mathbb{R}))$ such that each boundary representation is either hyperbolic or parabolic. Let $\mathcal{R}^{\prime}_{hp}(\Sigma)\subset\mathcal{R}_{hp}(\Sigma)$ be the subset such that $\phi|_{\pi_{1}(M_{i})}$ is non-abelian for each $i$. Hence $\mathcal{R}^{\prime}_{hp}(\Sigma)$ is open and dense in $\mathcal{R}_{hp}(\Sigma)$.
Lemma 4.7.
Let $\phi=\phi(a,b,c_{1},c_{2})=(A,B,C_{1},C_{2})\in\mathrm{Hom}(\pi_{1}(\Sigma_{0,4}),\mathrm{SL}(2,\mathbb{R}))$ be a representation such that $A,B$ are either parabolic or hyperbolic, see Figure 5. Assume that the restrictions $\phi|_{\pi_{1}(M_{i})}$, $i=1,2$, are non-abelian, where
Proof.
Since $\phi|_{\pi_{1}(M_{2})}$ is non-abelian, the element $C=C_{1}C_{2}$ must be either parabolic, elliptic, or hyperbolic. If $C$ is hyperbolic, then we may simply take $\phi^{\prime}=\phi$, which is the desired representation. If $C$ is elliptic, then by (2.7), we have $\mathrm{tr}([C_{2},C])\geq 2,$ with equality if and only if $C_{2}$ and $C$ commute. Since $\phi|_{\pi_{1}(M_{2})}$ is non-abelian, it follows that
$\mathrm{tr}([C_{2},C_{1}])=\mathrm{tr}([C_{2},C])>2,$
which implies that
(4.12) $(\mathrm{tr}(C_{1}),\mathrm{tr}(C_{2}),\mathrm{tr}(C))\in\Delta.$
If $C$ is parabolic, then by (2.6), we also have $\mathrm{tr}([C_{2},C])\geq 2$, with equality if and only if $C_{2}$ and $C$ commute or $C_{2}$ is hyperbolic. Again, since $\phi|_{\pi_{1}(M_{2})}$ is non-abelian, we conclude that (4.12) holds.
Now choose a path $(\mathrm{tr}(C_{1}),\mathrm{tr}(C_{2}),f_{t})$ in $\mathbb{R}^{3}$ such that $f_{t}$ is monotone with $f_{0}=\mathrm{tr}(C)$ and $|f_{1}|>2$. This is possible since we can choose $f_{1}$ to lie in the same connected component of $\Delta$ as $f_{0}$. By the path-lifting property of Theorem 4.6, there exists a path of representations
$\delta_{2,t}=(X_{t},Y_{t},(X_{t}Y_{t})^{-1})\in\mathrm{Hom}(\pi_{1}(M_{2}),\mathrm{SL}(2,\mathbb{R}))$
connecting $\delta_{2,0}=\phi|_{\pi_{1}(M_{2})}$ to $\delta_{2,1}$ such that $\chi(\delta_{2,t})=(\mathrm{tr}(C_{1}),\mathrm{tr}(C_{2}),f_{t})$. In particular, $\delta_{2,1}(c)$ is hyperbolic.
Since the entire path $\delta_{2,t}$ remains non-abelian, the matrices $\delta_{2,t}(c_{1})$ and $\delta_{2,t}(c_{2})$ are never $\pm I$. Because $\mathrm{tr}(X_{t})=\mathrm{tr}(C_{1})$ and $C_{1}=X_{0}$, there exists a path $U_{t}^{1}\subset\mathrm{SL}(2,\mathbb{R})$, with $U_{0}^{1}=I$, such that $\delta_{2,t}(c_{1})=X_{t}=U_{t}^{1}C_{1}(U_{t}^{1})^{-1}.$ Similarly, there exists $U_{t}^{2}\subset\mathrm{SL}(2,\mathbb{R})$ such that $\delta_{2,t}(c_{2})=Y_{t}=U_{t}^{2}C_{2}(U_{t}^{2})^{-1}.$
Next, we lift the restriction $\phi|_{\pi_{1}(M_{1})}$ along the path $\delta_{2,t}(c)^{-1}$ by solving
$\delta_{2,t}(c)=A_{t}B_{t},$
with $A_{t}$, $B_{t}$ parabolic. Define the path
$\alpha_{t}:=\left(\mathrm{tr}(A),\mathrm{tr}(B),\mathrm{tr}(\delta_{2,t}(c))\right)\in\mathbb{R}^{3}.$
Then
$\kappa(\alpha_{t})$ $=\mathrm{tr}(A)^{2}+\mathrm{tr}(B)^{2}+f_{t}^{2}-\mathrm{tr}(A)\mathrm{tr}(B)f_{t}-2$
$=\left(f_{t}-\tfrac{1}{2}\mathrm{tr}(A)\mathrm{tr}(B)\right)^{2}-\tfrac{1}{4}(\mathrm{tr}(A)^{2}-4)(\mathrm{tr}(B)^{2}-4)+2.$
If either $A$ or $B$ is hyperbolic, then $|\mathrm{tr}(A)|>2$ or $|\mathrm{tr}(B)|>2$, which implies that $\kappa(\alpha_{t})>2$, so $\alpha_{t}\in\Delta$ for all $t$. Then Theorem 4.6 guarantees the existence of a lift $\delta_{1,t}$ of $\phi|_{\pi_{1}(M_{1})}$ satisfying $\chi(\delta_{1,t})=\alpha_{t}$.
If both $A$ and $B$ are parabolic (i.e., $|\mathrm{tr}(A)|=|\mathrm{tr}(B)|=2$), then $\kappa(\alpha_{t})>2$ for all $t$ except possibly at a single point $t_{0}\in(0,1)$. Thus $\alpha_{t}\in\Delta$ for all $t\neq t_{0}$. Let $\delta_{1,t}^{\prime}$ be the lift of $\alpha_{t}$ on $0,t_{0})$ starting at $\phi|_{\pi_{1}(M_{1})}$, and define $\delta_{1,t_{0}}^{\prime}:=\lim_{t\nearrow t_{0}}\delta_{1,t}^{\prime}$. Since $\chi(\delta_{1,t_{0}}^{\prime})=(\pm 2,\pm 2,2)=\alpha_{t_{0}}$ and $f_{t}>2$ for $t>t_{0}$, we can lift $\alpha_{t}$ on $t_{0},1]$ to a path $\delta_{1,t}^{\prime\prime}$ starting from $\delta_{1,t_{0}}^{\prime}$. Gluing the segments, we obtain a continuous path
$\delta_{1,t}=(A_{t},B_{t},(A_{t}B_{t})^{-1})=(A_{t},B_{t},C_{t}),\quad 0\leq t\leq 1,$
with $\delta_{1,0}=\phi|_{\pi_{1}(M_{1})}$ and $\chi(\delta_{1,t})=\alpha_{t}$.
Since $\mathrm{tr}(C_{t}^{-1})=\mathrm{tr}(\delta_{2,t}(c))$ and $C_{0}^{-1}=\delta_{2,0}(c)$, there exists a path $\beta_{t}\subset\mathrm{SL}(2,\mathbb{R})$, with $\beta_{0}=I$, such that $C_{t}=\beta_{t}\delta_{2,t}(c)^{-1}\beta_{t}^{-1}.$ Then $\beta_{t}^{-1}\delta_{1,t}\beta_{t}$ is a lift of $\phi|_{\pi_{1}(M_{1})}$ along $\delta_{2,t}(c)^{-1}$.
Finally, glue the deformed restrictions $\delta_{1,t}$ and $\delta_{2,t}$ along $c$ to obtain a deformation $\delta_{t}$ of $\phi$, with $\phi^{\prime}:=\delta_{1}$ satisfying the conclusion. The conjugacy classes of $\delta_{t}(a)$, $\delta_{t}(b)$, $\delta_{t}(c_{1})$, and $\delta_{t}(c_{2})$ are preserved throughout. ∎
@@
The following lemma was originally established by Goldman in [8] for representations into $\mathrm{PSL}(2,\mathbb{R})$. Via the following map
$\pi_{*}:\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SL}(2,\mathbb{R}))\to\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{PSL}(2,\mathbb{R})),$
@@
which is a covering map on the inverse image of a connected component, see [8, Lemma 2.2]. Hence, any path in $\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{PSL}(2,\mathbb{R}))$ can be lifted to a path in $\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SL}(2,\mathbb{R}))$ starting from a fixed base point. Therefore, we obtain the following:
Lemma 4.8 (Goldman [8, Lemma 9.3]).
@@
We are now ready to prove the following key lemma, which generalizes [8, Lemma 10.1].
@@
Let $\Sigma=\Sigma_{g,n}$ be a surface of genus $g$ with $n\geq 1$ boundary components, and let $\Sigma=\bigcup_{i=1}^{|\chi(\Sigma)|}M_{i}$ be a maximal dual tree decomposition of $\Sigma$. Then any representation $\phi\in\mathcal{R}_{hp}(\Sigma)$ can be deformed to an interior hyperbolic representation $\phi^{\prime}\in\mathcal{R}_{hp}(\Sigma)$, i.e., a representation for which the restriction to each component $c\in\partial M_{i}\setminus\partial\Sigma$ is hyperbolic.
@@
We follow the strategy of [8, Lemma 10.1]. Since $\mathcal{R}^{\prime}_{hp}(\Sigma)$ is open and dense in $\mathcal{R}_{hp}(\Sigma)$, we may assume without loss of generality that $\phi\in\mathcal{R}^{\prime}_{hp}(\Sigma)$. We proceed by induction on the Euler characteristic $\chi(\Sigma)$. The base cases $\chi(\Sigma)=-1$ and $\chi(\Sigma)=-2$ are established by Lemmas 4.7 and 4.8.
@@
    $\partial M_{1}$ contains at least one component of $\partial\Sigma$;
@@
The remaining subsurfaces $\bigcup_{i>2}M_{i}$ consist of one or two connected components $N_{j}$. Let $c_{j}$ denote the boundary curve $c_{j}=\partial M_{2}\cap\partial N_{j}$, and let $c=\partial M_{1}\cap\partial M_{2}$; see Figure 5.

### v3

@@ 2.4. The case of general genus
If $\mathrm{tr}(C_{1})>2$, then we can similarly choose deformations $A_{1}(t),B_{1}(t)$ such that $\mathrm{tr}(C_{1}(t))$ decreases into the interval (2,2)(-2,2), while preserving the sign of $(C_{1}(t))_{21}$. In fact, by the path-lifting property [10, Corollary 7.8], we can construct deformations $A_{1}(t),B_{1}(t)$ such that $\mathrm{sgn}((C_{1}(t))_{21})$ remains constant throughout the deformation, where $C(t)$ is hyperbolic for $[0,1)$, and $C(1)$ is parabolic. Moreover, we can arrange that
@@
For representations satisfying $|\mathrm{T}(\phi)|<|\chi(\Sigma_{g,0})|$, the argument is the same as in Lemmas 2.20 and 2.23, as detailed in [10, Lemma 10.5]. This shows that for each such Toledo invariant value, the set of corresponding representations forms a connected component, and there are 2g32g-3 such values.
@@ 2.5. The case of $\mathrm{PSL}(2,\mathbb{R})$
Let $\pi:\mathrm{SL}(2,\mathbb{R})\to\mathrm{PSL}(2,\mathbb{R})$ denote the natural projection. This induces a map which is a covering map on its inverse of a connected component ([10, Lemma 2.2])
@@
Since $\phi^{\prime}|_{\pi_{1}(\Sigma_{g,0})}]$ and $\psi^{\prime}|_{\pi_{1}(\Sigma_{g,0})}]$ are connected in $\mathrm{Hom}(\pi_{1}(\Sigma_{g,0}),\mathrm{PSL}(2,\mathbb{R}))$ (see [10, Theorem A (i)]), and $\phi^{\prime}|_{\pi_{1}(\Sigma_{0,n})}$ and $\psi^{\prime}|_{\pi_{1}(\Sigma_{0,n})}$ are also connected, it follows that $\phi^{\prime}]$ and $\psi^{\prime}]$ are connected in $\mathrm{Ell}_{P}(\Sigma_{g,n})$. Consequently, $\phi]=[\mu(\phi)]$ and $\psi]$ lie in the same connected component of $\mathrm{Ell}_{P}(\Sigma_{g,n})$. ∎
@@ 3.2. The general case
Assume the lemma holds for all $\leq n$. For $n+1$, consider the maximal dual-tree decomposition as illustrated on the right in Figure 4. By [10, Lemma 10.1], we may assume that $D_{1},\dots,D_{n-2}$ are hyperbolic elements.
@@
By [10, Lemma 10.1], we may assume that $C_{-1}$ and $C_{1}$ are hyperbolic. If $\mathrm{T}(\phi_{1})<1$, then as in the case $n=3,g=0$, we may deform $\phi_{1}$ along a $\gamma$-path while fixing the conjugacy classes of $C_{-1}$ and $C_{0}$, so that the Toledo invariant of $\phi_{1}$ increases by 1.
@@
By [10, Lemma 10.1], we may assume that it is an interior hyperbolic representation with respect to a maximal dual tree decomposition, see also Lemma 4.11.
@@
Note that $\mathrm{Hyp}(\Sigma_{1,1})\cap\mathrm{T}^{-1}(0)$ is connected. This follows from the same argument as in [10, Lemma 10.5] and the sigma value of each representation in $\mathrm{Hyp}(\Sigma_{1,1})\cap\mathrm{T}^{-1}(0)$ is always equal to $1$ (see Table 2), since one can construct a representation $\phi_{1}\in\mathrm{Hyp}(\Sigma_{1,1})\cap\mathrm{T}^{-1}(0)$ such that $\phi_{1}(a_{1})$ or $\phi_{1}(b_{1})$ is elliptic. Furthermore, if the set $\mathrm{Hyp}(\Sigma_{0,g+n})\cap\mathrm{T}^{-1}(l)\cap\sigma^{-1}(a)$ is non-empty, by Theorem 3.3 and the fact that $\sigma^{-1}(a)$ is fixed, it is connected. By Theorem 3.3, any $\phi$ as above can be deformed into a representation of the form $\gamma\phi_{0}$, where $\gamma\in\Lambda_{g,n}$.
@@
Write $C_{3}:=\Pi_{i=1}^{g-1}\phi([a_{i},b_{i}])$ and $C_{2}:=\phi([a_{g},b_{g}])$, so that $C_{3}C_{2}C_{1}=I$. By [10, Lemma 10.1], we may assume that $C_{2}$ and $C_{3}$ are both hyperbolic. Since $\mathrm{T}(\phi)=\pm|\chi(\Sigma_{g,1})|=\pm(2g-1)$, and by the inductive assumption, we have $\sigma(C_{2})=\sigma(C_{3})=0$. Furthermore, the triple $(C_{1},C_{2},C_{3})$ forms a representation of $\Sigma_{0,3}$ with Toledo invariant zero if and only if $\sum_{i=1}^{3}\sigma(C_{i})$ is odd. Hence, $T((C_{1},C_{2},C_{3}))=\pm 1$ implies $\sigma(C_{1})=0$.
We now consider the case $n\geq 2$. Again, by [10, Lemma 10.1], we may assume that $C_{0}$, $D_{1},\dots,D_{n-2}$ are all hyperbolic (see Figure 4).
@@ 4.1. Pairs of pants and punctured torus
In this section, we consider the case of surfaces $\Sigma=\Sigma_{0,3}$ and $\Sigma_{1,1}$.
@@ 4.1.1. The case of pairs of pants
(4.2) $\sigma\colon\mathrm{Par}(\Sigma)\to\{\pm 1,\pm\sqrt{-1}\}^{n},\quad\phi\mapsto\left(\sigma(\phi(c_{1})),\dots,\sigma(\phi(c_{n}))\right).$
For any $\phi\in\mathrm{Par}(\Sigma)$, we typically use a vector $a\in\{\pm 1,\pm\sqrt{-1}\}^{n}$ to represent its sigma value, that is, $a=\sigma(\phi)$. For a given sigma value $a$, there is an associated vector, called the $s$-value, defined by
(4.3) $s=s_{a}=(s_{1},\dots,s_{n})\in\{\pm 1\}^{n},\quad\text{where}\quad s_{i}=\mathrm{Im}(a_{i})-\mathrm{Re}(a_{i}).$
Note that the $s$-value, like the Toledo invariant, satisfies the following property: if two representations in $\mathrm{Par}(\Sigma)$ are both liftings of the same representation in $\mathrm{Hom}(\pi_{1}(\Sigma_{g,n}),\mathrm{PSL}(2,\mathbb{R}))$, then their $s$-values are identical.
@@
(4.10) $\begin{split}\#\{\mathrm{Par}(\Sigma_{0,3})\}&=\#\{a_{1},a_{2}:\mu s=1,c\in\{0,2\}\}\\ &\quad+\#\{a_{1},a_{2}:\mu s=-1,c=0,s+\mu a^{2}\neq 0\}\\ &=2\,\#\{a_{1},a_{2}:\mu s=1\}+2\#\{a_{1},a_{2}:\mu s=-1\}\\ &=2\times 8+2\times 8=32.\end{split}$
@@ 4.2. Interior hyperbolic representations
In this section, we consider deforming a given representation to an interior hyperbolic representation with respect to a maximal dual tree decomposition. As an application, we examine the possible boundary sigma values for boundary parabolic representations with maximal Toledo invariant.
4.2.1. Interior hyperbolic representations
In [10, Lemma 10.1], Goldman proved that any boundary hyperbolic representation $\phi\in\mathrm{Hyp}(\Sigma)$ can be deformed to another boundary hyperbolic representation $\phi^{\prime}$, such that for a maximal dual tree decomposition $\Sigma=\bigcup_{i=1}^{|\chi(\Sigma)|}M_{i}$ (see [10, Section 3.8]), the restriction $\phi^{\prime}|_{\pi_{1}(M_{i})}$ is hyperbolic on $\partial M_{i}$ for each $i$. We refer to such a representation $\phi^{\prime}$ as an interior hyperbolic representation.
In this subsection, we demonstrate that any boundary parabolic representation can similarly be continuously deformed into one that is hyperbolic on every interior pair of pants for $g\geq 1$.
@@ 4.2.1. Interior hyperbolic representations
Theorem 4.6 (Goldman [10, Proposition 4.1, Theorem 4.3, Corollary 4.5]).
@@
Let $\Sigma=\Sigma_{g,n}$ be a surface of genus $g$ and $n$ boundaries. We assume that $\chi(\Sigma)\leq-1$. Let $\Sigma=\bigcup_{i=1}^{|\chi(\Sigma)|}M_{i}$ be a maximal dual tree decomposition. Let $\mathcal{R}_{h}(\Sigma)$ be the set of all representations $\phi\in\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SL}(2,\mathbb{R}))$ such that one of boundary representation is hyperbolic. Let $\mathcal{R}^{\prime}_{h}(\Sigma)\subset\mathcal{R}_{h}(\Sigma)$ be the subset such that $\phi|_{\pi_{1}(M_{i})}$ is non-abelian for each $i$. Hence, $\mathcal{R}^{\prime}_{h}(\Sigma)$ is open and dense in $\mathcal{R}_{h}(\Sigma)$, since there is at least one boundary component whose holonomy is hyperbolic and can be freely varied.
@@
The following lemmas were originally established by Goldman in [10] for representations into $\mathrm{PSL}(2,\mathbb{R})$. Via the following map
(4.14) $\pi_{*}:\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SL}(2,\mathbb{R}))\to\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{PSL}(2,\mathbb{R})),$
@@
which is a covering map on the inverse image of a connected component, see [10, Lemma 2.2]. Hence, any path in $\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{PSL}(2,\mathbb{R}))$ can be lifted to a path in $\mathrm{Hom}(\pi_{1}(\Sigma),\mathrm{SL}(2,\mathbb{R}))$ starting from a fixed base point. Therefore, we obtain the following:
Lemma 4.7 (Goldman [10, Lemma 9.3]).
@@
Lemma 4.8 (Goldman [10, Lemma 9.6]).
Let $\phi=\phi(a,b,c_{1},c_{2})=(A,B,C_{1},C_{2})\in\mathrm{Hom}(\pi_{1}(\Sigma_{0,4}),\mathrm{SL}(2,\mathbb{R}))$ be a representation such that one of $A,B,C_{1},C_{2}$ is hyperbolic, see Figure 5. Assume that the restrictions $\phi|_{\pi_{1}(M_{i})}$, $i=1,2$, are non-abelian, where
@@
We are now ready to prove the following key lemma, which generalizes [10, Lemma 10.1].
@@
Let $\Sigma=\Sigma_{g,n}$ be a surface of genus $g$ with $n\geq 1$ boundary components, and let $\Sigma=\bigcup_{i=1}^{|\chi(\Sigma)|}M_{i}$ be a maximal dual tree decomposition of $\Sigma$. Then any representation $\phi\in\mathcal{R}_{h}^{\prime}(\Sigma)$ can be deformed in $\mathcal{R}_{h}^{\prime}(\Sigma)$ to an interior hyperbolic representation $\phi^{\prime}\in\mathcal{R}_{h}^{\prime}(\Sigma)$, i.e., a representation for which the restriction to each component $c\in\partial M_{i}\setminus\partial\Sigma$ is hyperbolic, while boundary conjugacy classes are fixed.
@@
We follow the strategy of [10, Lemma 10.1]. We proceed by induction on the Euler characteristic $\chi(\Sigma)$. The base case $\chi(\Sigma)=-2$ is established by Lemmas 4.7 and 4.8.
@@
    $\partial M_{1}$ contains at least one component of $\partial\Sigma$, and the restriction of $\phi$ on this component is hyperbolic;
@@
The remaining subsurfaces $\bigcup_{i>2}M_{i}$ consist of at most three connected components $N_{j}$. Let $c_{j}$ denote the boundary curve $c_{j}=\partial M_{2}\cap\partial N_{j}$, and let $c=\partial M_{1}\cap\partial M_{2}$; see Figure 5.